OpenStudy (anonymous):

use the 5 steps for problem solving to answer the following question. Enrollment at a particular online university is 11,600. there are seven times as many students who work full time as there are students who do not. how many students work full time and how many do not.

OpenStudy (anonymous):

let x = number who work full time so no. who doi not = 11,600 - x seven times more who work s0 x = 7(11600 - x) solv this equation to get value of x and 11,600-x.

OpenStudy (anonymous):

8x = 7*11600 x = 10,150

OpenStudy (anonymous):

hey jimmy im tryimg to figure this problem

OpenStudy (anonymous):

ok - does what i wrote help?

OpenStudy (anonymous):

yes thx so much is x= 10,150 my answer

OpenStudy (anonymous):

yes - i got there in the end - i've got the most awful cold and my brain is a bit slow today

OpenStudy (anonymous):

oh bless tour heart jimmy.....hope u to feel better...i been in hospital i tore my rotator cuff in my shoulder lol

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

sounds painful! - i think i'd prefer my cold to that! you ok now?

OpenStudy (anonymous):

no doubt, but colds are just really horrible

OpenStudy (anonymous):

this virus is sweeping the swansea area - it mainly attacks the throat and it lasts about 2-3 weeks - i've had it about a week happy days! lol

OpenStudy (anonymous):

lol well i hope u feel better soon jimmy and thx for all your help =)

OpenStudy (anonymous):

thnx - yw

OpenStudy (anonymous):

=) an for the help with the other problem as well